A newspaper and a magazine cost £110 together. The newspaper cost £70 less than the magazine. Find the cost of each

Answer by vheroli(126)

Solution:
110-70=40
40/2=20
Answer:
Cost of newspaper= £20
Cost of magazine
= 20+70
= £90